# Theory Uncertainties in the Extraction of $\alpha_{s}$ from Drell-Yan at Small Transverse Momentum

### v1

@@ B.2 Approximate TNP orders
Figure -948: Left panel: Uncertainties in the $q_{T}$ spectrum at N3+0LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations.
In figure -953, we already included the results for $\alpha_{s}(m_{Z})$ at N3+0LL and N4+0LL. Here, we provide the corresponding plots for the $q_{T}$ spectrum and TNP pull plots. The Asimov fit follows the same setup as in section 4.2.3. The results at N3+0LL are shown in figure -948. In the left panel, we observe a significant reduction in uncertainty from the pre-fit to the post-fit bands. This reduction is particularly strong due to the high degree of correlation in the fit. Such a reduction may be suspicious, and indeed, the TNP pull plot in the right panel provides insights into the reasons for this behaviour. Some TNPs, particularly the beam functions, become strongly constrained by the data. When the constraints on these parameters become this tight, it suggests that the considered order may not be sufficient to correctly account for the remaining theory uncertainties. This is an example of overfitting the theory model, as previously discussed in section 4.2.3. Regarding the TNPs’ impacts on $\alpha_{s}(m_{Z})$, we find that $\gamma_{\nu}$, $S$ and $B_{qg}$ have the strongest impact on $\alpha_{s}$. This is consistent with what we observed in the N3+1LL case. In contrast, at N2+1LL, the pattern of TNP constraints and pulls was quite different.
Figure -947: Left panel: Uncertainties in the $q_{T}$ spectrum at N4+0LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations.
The results for the N4+0LL case are shown in figure -947. By construction, the central N4+0LL prediction is identical to that of N4LL. The reduction in uncertainty across the $q_{T}$ spectrum is less pronounced compared to the N3+0LL case, and similarly the constraints on the TNPs are no longer as tight as in the N3+0LL case.

### v2

@@ B.2 Approximate TNP orders
Figure 12: Left panel: Uncertainties in the $q_{T}$ spectrum at N3+0LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations.
In figure 7, we already included the results for $\alpha_{s}(m_{Z})$ at N3+0LL and N4+0LL. Here, we provide the corresponding plots for the $q_{T}$ spectrum and TNP pull plots. The Asimov fit follows the same setup as in section 4.2.3. The results at N3+0LL are shown in figure 12. In the left panel, we observe a significant reduction in uncertainty from the pre-fit to the post-fit bands. This reduction is particularly strong due to the high degree of correlation in the fit. Such a reduction may be suspicious, and indeed, the TNP pull plot in the right panel provides insights into the reasons for this behaviour. Some TNPs, particularly the beam functions, become strongly constrained by the data. When the constraints on these parameters become this tight, it suggests that the considered order may not be sufficient to correctly account for the remaining theory uncertainties. This is an example of overfitting the theory model, as previously discussed in section 4.2.3. Regarding the TNPs’ impacts on $\alpha_{s}(m_{Z})$, we find that $\gamma_{\nu}$, $S$ and $B_{qg}$ have the strongest impact on $\alpha_{s}$. This is consistent with what we observed in the N3+1LL case. In contrast, at N2+1LL, the pattern of TNP constraints and pulls was quite different.
Figure 13: Left panel: Uncertainties in the $q_{T}$ spectrum at N4+0LL relative to the N4LL, before (yellow band) and after (orange band) profiling the TNPs. The different lines show the post-fit relative impact of each TNP. Right panel: Post-fit constraints on the TNPs (error bars) and their impact on $\alpha_{s}(m_{Z})$, with the solid (dashed) grey band showing the impact of the post-fit downward (upward) TNP variations.
The results for the N4+0LL case are shown in figure 13. By construction, the central N4+0LL prediction is identical to that of N4LL. The reduction in uncertainty across the $q_{T}$ spectrum is less pronounced compared to the N3+0LL case, and similarly the constraints on the TNPs are no longer as tight as in the N3+0LL case.